OpenStudy (anonymous):

√x+6 The remains of a structure on an archeological site are laid out in the shape of a rectangle. The area of the rectangle is 12 square meters. The width is 3 meters and the length is as shown in this statement. What is the value of x? a) 10 b) 4 c) 12 d) 16

OpenStudy (anonymous):

Did you mean root(x) + 6? Or did you mean root(x+6)?

OpenStudy (mathmate):

3*length = 12 therefore length = 12/3=4 m Solve 4=sqrt(x+6) to find x.